OpenStudy (anonymous):

If you figure this out, you're absolutely amazing because it's irritating me: A jet travels at 402 m/s. How long will it take to travel 11.3 inches?

OpenStudy (anonymous):

Well, this is a case of metric to US conversion. If you remember, a centimeter (cm) is 1/2.54 inches. So 1 inch = 2.54 centimeters. There are 100 cm in a meter. So there are .0254 m = 1 inch. 402 m/s would translate to 402 m/s * 1 inch/.0254m. Do this conversion and then you can find the answer easily. Thanks, Sean Walsh Facebook: Tutor Sean http://www.tutorsean.net

OpenStudy (anonymous):

Since 11.3 inches is = 0.28702m, Then,0.28702m/(402 m/s) = 0.0007139801 seconds

OpenStudy (anonymous):

Nicely put.

OpenStudy (anonymous):

i could have sworn i just did this problem in a previous post...

OpenStudy (anonymous):

Anytime

OpenStudy (radar):

1 meter is approx 39.37 in. So the jet is traveling 402 x 39.37 in. per sec or 15826.74 in/sec. 11.3/15826.74=.0007392 sec t=(d/r) 739.2 microseconds

OpenStudy (anonymous):

You are right. 1 meter is approx. 3 ft. It's been a long weekend. ;)